OpenStudy (anonymous):

A test consisted of 20 multiple choice questions plus 10 true or false questions. Barry got 4 questions wrong, what percent of his answers were correct.. I know the answer is 87% but how do you get that?

OpenStudy (bahrom7893):

if he got four wrong, that means he had 16 correct, so if you have a total of 20 questions: 20 - 100% 16 - x% x = (16*100/20)%

OpenStudy (bahrom7893):

Oh wait nevermind, there were 30 questions in total

OpenStudy (bahrom7893):

So it's not 20 - 100% 16 - x% x = (16*100/20)%, it's: 30 - 100% 26 - x% x = (26*100/30)% = 86.6666667%

OpenStudy (bahrom7893):

which is about 87%

OpenStudy (anonymous):

thanks:)